[Hallc_running] [New Logentry] Hall C status update (Dec. 19)

gaskelld at jlab.org gaskelld at jlab.org
Tue Dec 19 09:40:04 EST 2017

Logentry Text:
Beam returned late swing shift yesterday after set up of the 5th pass RF separation to Halls B and D. After checking the beam position with the carbon hole target, time was dedicated to checking and optimizing the SHMS tune with carbon elastics at 7.5 degrees. In the end, Mark and Holly settled on slightly different Q1/Q2/Q3 settings than had been determined from the larger angle (10 degree) SHMS data.

Today, we are condensing the hydrogen target which requires the hall to be in Controlled Access. After the cooldown, we will perform a delta scan in the SHMS with the new tune.

Also - at the 8 am meeting today, it was decided that the run will stop at 8 am on Thursday - no shift workers are needed for the Thursday day shift.

There will be an RC meeting today at 1 pm in the Counting House meeting room (remote connection information on the wiki).


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3506900
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20171219/b95300cd/attachment.html>

More information about the Hallc_running mailing list